OpenStudy (anonymous):

calculus question.... let f(x) = (x^2+4)/x write an equation of the normal line to the graph of f at the point where x=2

OpenStudy (anonymous):

\(f(2)=4\) so the point is \((2,4)\) now you need the slope did you find the derivative?

OpenStudy (anonymous):

i suppose you could use the quotient rule, but it is easier to divide and get \[f(x)=x+\frac{4}{x}\] so \(f'(x)=1-\frac{4}{x^2}\) and \[f'(2)=1-1=0\] hmm is this a trick question?

OpenStudy (anonymous):

no matter, tangent line is horizontal, normal line is vertical, and vertical line through \((2,4)\) is \(x=2\)

OpenStudy (anonymous):

i got as the derivative of the f(x) x-4

OpenStudy (anonymous):

it is \[f(x)=\frac{x^2+4}{x}\]?

OpenStudy (anonymous):

yes so the derivative is x-4 and i got the equation for the tangent line at x=1 but unsure how to find the normal line at x=2

OpenStudy (anonymous):

hold on the derivative is not \(x-4\)

OpenStudy (anonymous):

\(x-4\) is the derivative of \(\frac{1}{2}x^2-4x\) not \(\frac{x^2+4}{x}\)

OpenStudy (anonymous):

i am not sure what rule you used, but you can either use the quotient rule, or more easily divide and take the derivative piece by piece

OpenStudy (anonymous):

i used quotient rule.... \[gf'-fg'/g^2\]

OpenStudy (anonymous):

i got y-4=-1/3(x-2)

OpenStudy (anonymous):

ok lets get the derivative correct first, then we can find the slope

OpenStudy (anonymous):

\[\frac{gf'-fg'}{g^2}\] with \[f(x)=x^2+4,f'(x)=2x,g(x)=x,g'(x)=1\] you get \[\frac{x\times 2x-(x^2+4)}{x^2}=\frac{x^2+4}{x^2}\]

OpenStudy (anonymous):

damn typo it is \[\frac{x^2-4}{x^2}\] sorry

OpenStudy (anonymous):

its a test review.... :( im gunna fail lol

OpenStudy (anonymous):

although it was easier to compute the derivative of \[f(x)=x+\frac{4}{x}\] then \[f'(x)=1-\frac{4}{x^2}\]

OpenStudy (anonymous):

nah, just some practice, but look the only way you are going to get a derivatie that is a nice line line \(x-4\) is if you start with a parabola like \(\frac{1}{2}x^2-4x\)

OpenStudy (anonymous):

if you have a quotient expect to get a numerator and a denominator for the derivative the derivative of a rational function will be another rational function, not a line

OpenStudy (anonymous):

mm i got it its 2x-4 cause the top x2 and bottom x2 cancel out.

OpenStudy (anonymous):

whoa nelly stop cancelling!

OpenStudy (anonymous):

don't expect to cancel ever. there is rarely cancelling in the quotient rule, and in any case \(x^2\) is not a factor of the numerator, careful with the algebra, that will mess you up for sure

OpenStudy (anonymous):

but don't forget, your job is to evaluate the derivative, so no need to "simplify" and make a mistake just plug in 2 and see what you get

OpenStudy (anonymous):

its the only way i see it working lol the tops neg and bottom is positive... because f' is 2 and g' is 1.

OpenStudy (anonymous):

\(f'(x)=2x\) not \(2\)

OpenStudy (anonymous):

ok hold on. i think i got. thanks :)

OpenStudy (anonymous):

if you are taking a test, remember that the derivative of a rational function is another rational function, not a line or any other polynomial only "combine like terms" in the numerator, do not cancel anything

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ok so wait i got the derivative. so back to original question how do you solve for the slope. and new question how do you get the points where f has a horizontal tangent?